OpenStudy (anonymous):

sqrt x+33 plus sqrt x-2 =7

OpenStudy (anonymous):

\[\sqrt{x+33} + \sqrt{x-2} = 7\] Is that what you're trying to say?

OpenStudy (anonymous):

Please reply, or I can't help you.

OpenStudy (anonymous):

I'd answered htis before.... Considering you mean \[\sqrt{x+33}+\sqrt{x-2}=7\] You multiply by \[\sqrt{x+33}-\sqrt{x-2}\] And you get \[(x+33)-(x-2)=35=7(\sqrt{x+33}-\sqrt{x-2})\leftarrow \rightarrow \sqrt{x+33}-\sqrt{x-2}=5\] So you have a system \[\sqrt{x+33}+\sqrt{x-2}=7 \] \[\sqrt{x+33}-\sqrt{x-2}=5\] You add them up and get \[2\sqrt{x+33}=12\] \[\sqrt{x+33}=6\] Meaning \[x+33=36\leftarrow \rightarrow x=3\]